OpenStudy (anonymous):

Simple chemistry formula that I can't seem to figure out how they got T1 by itself.... v1/ t1 = v2 / t2 t1 = (v1/v2) t2 how??

jimthompson5910 (jim_thompson5910):

v1/t1 = v2/t2 v1*t2 = t1*v2 ... cross multiply t1*v2 = v1*t2 ... flip the equation (we want t1 on the left side) t1 = (v1*t2)/(v2) ... Divide both sides by v2. t1 = (v1/v2)*t2 ... rearrange terms

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw